OpenStudy (anonymous):

calculates the limit l'hopital rule with \[\lim_{x \rightarrow 0 } a sen(x) - x e ^{x}/x ^{^{2}}\]

OpenStudy (anonymous):

need help....

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

@amistre64 @akash123

OpenStudy (anonymous):

wat's the 1st term?

OpenStudy (anonymous):

arc sec(x)

OpenStudy (anonymous):

can u write down the question again?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} a sen (x) e ^{x} / x ^{2}\]

OpenStudy (anonymous):

not is arc sen is a sen

OpenStudy (anonymous):

I am not aware of function ...a sen (x)

OpenStudy (anonymous):

hmmm... sorry is a sin (x)

OpenStudy (anonymous):

then differentiate numerator and denominator..

OpenStudy (anonymous):

i dont understand how to solve this question with rule of hopital

OpenStudy (anonymous):

if u have to find the limit of type 0/0 0r infinity /infinity then lim x--->a f(x)/g(x) = lim x--->a f'(x)/g'(x) either f(a) n g(a) both tend to 0 or infinity

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

lim x--->a f(x)/g(x) = lim x--->a f'(x)/g'(x)...this's called l'hospital's rule either f(a) n g(a) both tend to 0 or infinity

OpenStudy (anonymous):

is it fine?

OpenStudy (anonymous):

e.g lim x--->0 asinx (e^x)/x^2 ( form of 0/0) ...right?

OpenStudy (anonymous):

lim x--->0 asinx (e^x)/x^2 = lim x--->0 a[ e^x sinx + e^x cosx]/ 2x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

have u written down the question correctly?

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

lim x--->0 asinx (e^x)/x^2 ...is this the right question?

OpenStudy (anonymous):

if yes then lim x--->0 asinx (e^x)/x^2 = lim x--->0 a[ e^x sinx + e^x cosx]/ 2x = lim x--->0 a/ 2x and limit does not exist

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u know the answer?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

\[\Large \lim_{x \rightarrow 0} \frac{asin(x)e^{x}}{x^2}\] is this is your question ?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

limit doesn't exist for this question

OpenStudy (anonymous):

might be ...it should be sin^2x instead of sinx

OpenStudy (anonymous):

but now u know the l'hospital rule...u can try other questions

OpenStudy (anonymous):

ok by applying the limits you get 0/0 form which means this question is condiate for L Hopital rule in this rule you differentiate both numerator and denominator by taking derivative of numerator we have by product rule of differentiation \[\Large \frac{d}{dx}(asin(x)*e^x)=a \frac{d}{dx}(\sin(x)*e^x)=a[\cos(x)e^x+\sin(x)e^x)]\] the derivative of denominator is 2x so \[\Large \lim_{x \rightarrow 0} \frac{a[\cos(x)e^x+\sin(x)e^x]}{2x}\] now apply the limit \[\Large \frac{a[\cos(0)e^0+\sin(0)e^(0)]}{2(0)}\] where \[\Large \cos(0)=1\] \[\Large \sin(0)=0\] \[\Large e^{0}=1\] let me know what u get after the simplification of above ..

OpenStudy (anonymous):

1/0

OpenStudy (anonymous):

you missed a it is a/0 so it means \[\Large \frac{a}{0}=\infty\] should be the answer .

OpenStudy (anonymous):

yes...thanks :)))

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

Sir:)